Tom Scott is the owner, president, and primary salesperson for Scott Manufacturing. Because of this, the company's profits are driven by the amount of work Tom does. If he works 40 hours each week, the company's EBIT will be $630,000 per year; if he works a 50-hour week, the company's EBIT will be $785,000 per year. The company is currently worth $4.00 million. The company needs a cash infusion of $2.10 million, and it can issue equity or issue debt with an interest rate of 10 percent. Assume there are no corporate taxes.

a. What are the cash flows to Tom under each scenario? (Enter your answers in dollars, not millions of dollars, e.g. 1,234,567. Do not round intermediate calculations.)

Scenario-1
Debt issue:

Cash flows
40-hour week $
50-hour week $


Scenario-2
Equity issue:

Cash flows
40-hour week $
50-hour week $


b. Under which form of financing is Tom likely to work harder?

  • Equity issue

  • Debt issue
Answer #1

Debt Issue

40 hour week:

EBIT = 630,000

Interest for a year = Funds x interest rate = 2,100,000 x 10% = 210,000

Hence EBT = EBIT - Interest = 630,000 - 210,000 = 420,000

No taxes. Hence, cashflows = EBT = 420,000

50 hour week:

EBIT = 785,000

Interest for a year = Funds x interest rate = 2,100,000 x 10% = 210,000

Hence EBT = EBIT - Interest = 785,000 - 210,000 = 575,000

No taxes. Hence, cashflows = EBT = 575,000

Please fill your table like this:

Scenario-1
Debt issue:

Cash flows
40-hour week $ 420,000
50-hour week $ 575,000

Equity issue:

There will be no interest.

There are no taxes. Hence cash flows = EBIT

However due to equity raise, shareholding of the existing shareholder Tom will be diluted.

Pre equity raise valuation = 4 mn

Equity raise = 2.1 mn

Post equity raise valuation = 4 + 2.1 = 6.1 mn

%age shareholding of Tom now = Pre equity raise valuation / Post equity raise valuation = 4/6.1 = 65.57%

Hence, Tom will now have access to only 65.57% of the free cash flows to equity holders.

Under 40 hours week situation:

Cash flows to Tom = 65.57% x 630,000 = $ 413,115

Under 50 hours week situation:

Cash flows to Tom = 65.57% x 785,000 = $ 514,754

Please fill the table as shown below:

Scenario-2
Equity issue:

Cash flows
40-hour week $ 413,115
50-hour week $ 514,754

b. Under which form of financing is Tom likely to work harder?

  • Equity issue

  • Debt issue

Since cash flows for Tom is more in case of debt issue, Tom is likely to work harder under Debt Issue.